OpenStudy (1645323):

SOlve by Factoring:6x^4-5x^2-4=0

OpenStudy (1645323):

@phi

OpenStudy (phi):

you did not finish the other problem. this one is complicated.

OpenStudy (1645323):

I got the other problem

OpenStudy (1645323):

@phi

OpenStudy (phi):

I'll type out the final answer in the other post.

OpenStudy (phi):

in this problem \[ 6x^4-5x^2-4=0 \] you list the 6 (from the x^4) and the -4 (last term) 6* -4 = -24 the minus means the factors are different signs list the factors that give 24 1,24 2,12 3,8 4,6 look for a pair that add up to -5 when they have different signs (the bigger number must be negative) 1-24= -23 2 - 12 = -10 3 - 8 = -5 <---- this one works 4-6= -2

OpenStudy (phi):

the answer will be (2x^2 ) (3 x^2 ) we need a -8 which means we want (2x^2 )( 3x^2 - 4) the other number must be +1 (2x^2+1 )( 3x^2 - 4) This is a hard problem to explain clearly.

OpenStudy (1645323):

ok thanks